# An Extended Modified Kadomtsov-Petviashvili Equation: Ermakov-Painlevé II Symmetry Reduction with Moving Boundary Application

### 17 Mar 2026 v1

@@ Abstract
1 School of Mathematics and Statistics,
University of New South Wales, Sydney, Australia.
Email: c.rogers@unsw.edu.au
2 Depto. de Matemática, Facultad de Ciencias Exactas y Naturales,
Universidad de Buenos Aires & IMAS-CONICET.
Ciudad Universitaria. Pabellón I (1428), Buenos Aires, Argentina
Email: pamster@dm.uba.ar
@@ 1 Introduction
Hybrid Ermakov-Painlevé II systems were originally derived in [1] via wave packet representations admitted by multi-dimensional coupled nonlinear Schrö-dinger systems incorporating de Broglie-Bohm potential terms. Therein, in particular, the canonical single component Ermakov-Painlevé II equation was shown to occur notably in the analysis of transverse wave propagation in a generalised Mooney-Rivlin hyperelastic material. It has subsequently been derived in such diverse physical applications as cold plasma physics [2], Korteweg capillarity theory [3] and in the analysis of Dirichlet boundary value problems which arise out of the classical Nernst-Planck electrolytic system [4]. Underlying integrable structure in multi-component Ermakov-Painlevé II systems has been delimited in [5] wherein an admitted Ermakov system invariant was applied to construct an algorithmic solution procedure involving a Painlevé II connection. A Bäcklund transformation admitted by Painlevé II was applied iteratively therein and the important link between the Ermakov-Painlevé II and integrable Painlevé XXXIV equation set down.
@@
A novel variant with temporal modulation of the modified Kadomtsev–Petviashvili equation of 2+1-dimensional soliton theory is presented here which admits Ermakov-Painlevé II symmetry reduction. The latter is applied to obtain exact solution to a class of nonlinear moving boundary problems of Stefan-type associated. A 2+1-dimensional generalisation of involutory-type transformations with origin in autonomisation of the canonical Ermakov-Ray-Reid system in [29] is applied to embed the extended mKP equation in a wide associated class with temporal modulation which inherits admittance of Ermakov-Painlevé II sysmmetry reduction.
@@ 2 Ermakov-Painlevé II Symmetry Reduction
$\begin{array}[]{l}m(t+a)^{m-1}\Psi+(t+a)^{m-1}(-n\xi)\Psi^{\prime}+(t+a)^{m-3n}\Psi^{{}^{\prime\prime\prime}}\\[8.53581pt] \qquad-3(t+a)^{3m-n}\Psi^{2}\Psi^{\prime}+\lambda(t+a)^{\mu-3m-n}\Psi^{-4}\Psi^{\prime}=0,\end{array}$ (6)
@@
$\begin{array}[]{l}m\Psi-(n\xi\Psi)^{\prime}+n\Psi+(t+a)^{1-3n}\Psi^{{}^{\prime\prime\prime}}\\[8.53581pt] \qquad-3(t+a)^{2m+1-n}\Psi^{2}\Psi^{\prime}\\[8.53581pt] \qquad+\lambda(t+a)^{\mu-4m-n+1}\Psi^{-4}\Psi^{\prime}=0\end{array}$ (7)
@@
$\Psi^{{}^{\prime\prime\prime}}-3\Psi^{2}\Psi^{\prime}-(1/3)(\xi\Psi)^{\prime}+\lambda\Psi^{-4}\Psi^{\prime}=0\ .$ (8)
@@
$\Psi^{{}^{\prime\prime}}-\Psi^{3}-(1/3)\xi\Psi-(\lambda/3)\Psi^{-3}=\zeta\ ,\ \zeta\ \epsilon\ \mathbb{R}$ (9)
which, on appropriate scalings $\Psi=\gamma w^{*}$, and $\xi=\epsilon z$ and with $\zeta=0$ reduces to the canonical Ermakov-Painlevé II equation [1]
$w^{*}=2w^{*3}+2w^{*}+\chi w^{*-3}$ (10)
with $\chi=(\lambda/3)\gamma^{-4}\epsilon^{2}$.
@@ 3 A Class of Stefan-Type Moving Boundary Problems
I. $U_{xx}-U^{3}-(\lambda/3)(t+a)^{-2}U^{-3}=L_{m}S^{i}\dot{S}$
on $x+\alpha^{*}y=S(t)=\gamma(t+a)^{1/3},\,t>0$.
@@ I. $U_{xx}-U^{3}-(\lambda/3)(t+a)^{-2}U^{-3}=L_{m}S^{i}\dot{S}$
$\Psi^{{}^{\prime\prime}}(\gamma)-\Psi^{3}(\gamma)-(\lambda/3)\Psi^{-3}(\gamma)=L_{m}(t+a)S^{i}\dot{S}\ .$ (12)
@@
$\Psi^{{}^{\prime\prime}}(0)-\Psi^{3}(0)-(\lambda/3)\Psi^{-3}(0)=H_{0}(t+a)^{k+1}$ (15)
@@ 4 An Airy Reduction
$\rho_{zz}=(\rho_{z})^{2}/2\rho+2\rho^{2}+z\rho+2\delta/\rho\ ,\quad\delta\ \epsilon\ \mathbb{R}$ (16)
which is equivalent to the classical integrable Painlevé XXXIV equation (qv [3], [5]). In particular, it admits a particular solution
@@
wherein $w(z)$ is governed by the canonical Painlevé II equation. Thus, the Ermakov-Painlevé II equation in turn admits a class of exact solutions
@@
It is remarked that the iterative action of the Bäcklund transformation admitted by Painlevé II as applied in the analysis of boundary problems for the Nernst-Planck system in [31] may be used to generate with (21) as seed solution a wide class of associated solutions $\Psi$ of the Ermakov-Painlevé II equation and thereby of moving boundary problems of which the latter constitutes a symmetry reduction.
@@ 5 Temporal Modulation via a Class of Involutory Transformations
$\begin{array}[]{c}dt^{**}=\rho^{*-2}dt^{*}=dt\ ,\quad dx^{**}=dx\ ,\quad dy^{**}=dy\ ,\\[8.53581pt] U^{**}=U^{*}/\rho^{*}=U\ ,\quad\rho^{**}=1/\rho^{*}=\rho\end{array}$ (26)
@@
$\rho^{*}_{t^{*}t^{*}}+\omega(t^{*})\rho^{*}=\mathcal{E}/\rho^{*3}\ ,\quad\mathcal{E}\ \epsilon\ \mathbb{R}\ .$ (27)

### 30 Mar 2026 v2

# Open Communications in Nonlinear Mathematical Physics ]ocnmp[ Vol.6 (2026) pp AbstractReferences Article

@@ 1 Introduction
Hybrid Ermakov-Painlevé II systems were originally derived in [1] via wave packet representations admitted by multi-dimensional coupled nonlinear Schrödinger systems incorporating de Broglie-Bohm potential terms. Therein, in particular, the canonical single component Ermakov-Painlevé II equation was shown to occur notably in the analysis of transverse wave propagation in a generalised Mooney-Rivlin hyperelastic material. It has subsequently been derived in such diverse physical applications as cold plasma physics [2], Korteweg capillarity theory [3] and in the analysis of Dirichlet boundary value problems which arise out of the classical Nernst-Planck electrolytic system [4]. Underlying integrable structure in multi-component Ermakov-Painlevé II systems has been delimited in [5] wherein an admitted Ermakov system invariant was applied to construct an algorithmic solution procedure involving a Painlevé II connection. A Bäcklund transformation admitted by Painlevé II was applied iteratively therein and the important link between the Ermakov-Painlevé II and integrable Painlevé XXXIV equation set down.
@@
A novel variant with temporal modulation of the modified Kadomtsev–Petviashvili equation of 2+1-dimensional soliton theory is presented here which admits Ermakov-Painlevé II symmetry reduction. The latter is applied to obtain exact solution to a class of nonlinear moving boundary problems of Stefan-type associated. A 2+1-dimensional generalisation of involutory-type transformations with origin in autonomisation of the canonical Ermakov-Ray-Reid system in [29] is applied to embed the extended mKP equation in a wide associated class with temporal modulation which inherits admittance of Ermakov-Painlevé II symmetry reduction.
@@ 2 Ermakov-Painlevé II Symmetry Reduction
$\begin{array}[]{l}m(t+a)^{m-1}\Psi+(t+a)^{m-1}(-n\xi)\Psi^{\prime}+(t+a)^{m-3n}\Psi{{}^{\prime\prime\prime}}\\[8.53581pt] \qquad-3(t+a)^{3m-n}\Psi^{2}\Psi^{\prime}+\lambda(t+a)^{\mu-3m-n}\Psi^{-4}\Psi^{\prime}=0,\end{array}$ (6)
@@
$\begin{array}[]{l}m\Psi-(n\xi\Psi)^{\prime}+n\Psi+(t+a)^{1-3n}\Psi{{}^{\prime\prime\prime}}-3(t+a)^{2m+1-n}\Psi^{2}\Psi^{\prime}\\[8.53581pt] \qquad+\lambda(t+a)^{\mu-4m-n+1}\Psi^{-4}\Psi^{\prime}=0\end{array}$ (7)
@@
$\Psi{{}^{\prime\prime\prime}}-3\Psi^{2}\Psi^{\prime}-(1/3)(\xi\Psi)^{\prime}+\lambda\Psi^{-4}\Psi^{\prime}=0.$ (8)
@@
$\Psi{{}^{\prime\prime}}-\Psi^{3}-(1/3)\xi\Psi-(\lambda/3)\Psi^{-3}=\zeta,\ \zeta\ \in\mathbb{R}$ (9)
which, on appropriate scalings $\Psi=\gamma w^{*}$, and $\xi=\epsilon z$ and with $\zeta=0$ reduces to the canonical Ermakov-Painlevé II equation [1] for $w^{*}(z)$:
$(w^{*})^{\prime\prime}=2w^{*3}+zw^{*}+\chi w^{*-3}$ (10)
with $\chi=(\lambda/3)\gamma^{-4}\epsilon^{2}$ upon choosing the scaling parameters so that ${e^{3}}/3=1$ and $\epsilon^{2}\gamma^{2}=2$, namely $\epsilon=3^{1/3}$, $\gamma={2^{1/2}}/\epsilon$.
@@ 3 A Class of Stefan-Type Moving Boundary Problems
I. $U_{xx}-U^{3}-(\lambda/3)(t+a)^{-2}U^{-3}=L_{m}S^{i}\dot{S}$ on $x+\alpha^{*}y=S(t)=\gamma(t+a)^{1/3},\,t>0$.
@@
$\Psi{{}^{\prime\prime}}(\gamma)-\Psi^{3}(\gamma)-(\lambda/3)\Psi^{-3}(\gamma)=L_{m}(t+a)S^{i}\dot{S}.$ (12)
@@
$\Psi{{}^{\prime\prime}}(0)-\Psi^{3}(0)-(\lambda/3)\Psi^{-3}(0)=H_{0}(t+a)^{k+1}$ (15)
@@ 4 An Airy Reduction
$\rho_{zz}=(\rho_{z})^{2}/2\rho+2\rho^{2}+z\rho+2\delta/\rho\ ,\quad\delta\in\mathbb{R}$ (16)
which is equivalent to the classical integrable Painlevé XXXIV equation (qv [3], [5]). In particular, it admits the solution
@@
wherein $w(z)$ is governed by the canonical Painlevé II equation [3]. Thus, the Ermakov-Painlevé II equation in turn admits a class of exact solutions
@@
It is remarked that the iterative action of the Bäcklund transformation admitted by Painlevé II as applied in the analysis of boundary problems for the Nernst-Planck system in [31] may be used to generate, with (21) as seed solution, a wide class of associated solutions $\Psi$ of the Ermakov-Painlevé II equation and thereby of moving boundary problems of which the latter constitutes a symmetry reduction.
@@ 5 Temporal Modulation via a Class of Involutory Transformations
$\begin{array}[]{c}dt^{**}=\rho^{*-2}dt^{*}=dt\ ,\quad dx^{**}=dx\ ,\quad dy^{**}=dy,\\[8.53581pt] U^{**}=U^{*}/\rho^{*}=U\ ,\quad\rho^{**}=1/\rho^{*}=\rho,\end{array}$ (26)
@@
$\rho^{*}_{t^{*}t^{*}}+\omega(t^{*})\rho^{*}=\mathcal{E}/\rho^{*3}\ ,\quad\mathcal{E}\ \in\mathbb{R}.$ (27)
@@ 6 Conclusion
Acknowledgements
We thank the anonymous reviewer for the careful reading of the manuscript and his/her fruitful suggestions.